OpenStudy (anonymous):

find the value of a such that the equation (x-a)(x-12)+2 can be written in the form of (x-b)(x-c)

OpenStudy (anonymous):

Multiplying it out and grouping it into the usual ax^2 + bx + c form would probably be a good start.

OpenStudy (anonymous):

acc it doesnt.......

OpenStudy (anonymous):

acc it does....

OpenStudy (anonymous):

i tried it doesnt...coz it comes out to b x^2-(a+12)+12a+2 and fr this equation to hve 2 distinct roots discriminant of this equation(D) =>0 bt i wsnt able to get value of a frm tht.......

OpenStudy (anonymous):

(x-a)(x-12)+2 = x^2 -ax -12x -12a + 2 = x^2 - x(12+a) - (12a-2) where your b = (12+a) and your c = (12a-2)

OpenStudy (anonymous):

it shuld be +12a+2

OpenStudy (anonymous):

Sorry, typo...point is, u now have a quadratic which u can solve.

OpenStudy (anonymous):

no try it.......if u hve...plz do tel me the ans.....

OpenStudy (anonymous):

Well, it is what it is, the roots are : http://www.wolframalpha.com/input/?i=x^2+-+x+%2812%2Ba%29+%2B+%2812a+%2B2%29+%3D+0+for+x